# On a mean-field Pontryagin minimum principle for stochastic optimal control

### v1

@@ Abstract.
This papers outlines a novel extension of the classical Pontryagin minimum (maximum) principle to stochastic optimal control problems. Contrary to the well-known stochastic Pontryagin minimum principle involving forward-backward stochastic differential equations, the proposed formulation is deterministic and of mean-field type. The Hamiltonian structure of the proposed Pontryagin minimum principle is achieved via the introduction of an appropriate gauge variable. The gauge freedom can be used to decouple the forward and reverse time equations; hence simplifying the solution of the underlying boundary value problem. We also consider infinite horizon discounted cost optimal control problems. In this case, the mean-field formulation allows converting the computation of the desired optimal control law into solving a pair of forward mean-field ordinary differential equations. The proposed mean-field formulation of the Pontryagin minimum principle is tested numerically for a controlled inverted pendulum and a controlled Lorenz-63 system.
@@ 1. Introduction
Here $B_{t}$ denotes $d_{x}$-dimensional Brownian motion, $\Sigma\in\mathbb{R}^{d_{x}\times d_{x}}$ the symmetric positive definite diffusion matrix, $R\in\mathbb{R}^{d_{u}\times d_{u}}$ a symmetric positive definite weight matrix, $G(x)\in\mathbb{R}^{d_{x}\times d_{u}}$ the possibly position-dependent control matrix, $c(x)\geq 0$ the running cost, $f(x)\geq 0$ the terminal cost, and $\gamma>0$ the discount factor. See, for example, [2] for more details. We also introduce the weighted norm $\|\cdot\|_{R}$ via
@@
The aim is to find the closed loop control law $U_{t}=u_{t}(X_{t})$ that minimizes the cost $J_{T}$ over the set of admissible control laws.
@@
It is well known, that the associated optimal control problems for ordinary differential equations (ODEs); i.e. $\Sigma=0$ in (1), can be solved by the Pontryagin minimum (maximum) principle (PMP) [21]. The PMP has been extended to SDEs (1) in terms of forward-backward SDEs [20, 2]. In this paper, we provide an alternative extension via a deterministic mean-field formulation, which leads to interacting particle implementations in the form of Hamiltonian ODEs. The formulation is attractive from a computational perspective since it allows for a decoupling of the forward and backward evolution equations for finite horizon optimal control problems and a forward simulation for infinite horizon optimal control problems. The proposed mean-field approach also delivers closed loop control laws in contrast to the classical PMP for ODEs [21]. The proposed methodology can then also be used to approximate viscosity solutions for controlled ODEs by setting $\Sigma=\epsilon I$ for sufficiently small $\epsilon$ [3].
@@
Mean-field formulations and their interacting particle approximations have been considered before in the context of the Hamilton–Jacobi–Bellman (HJB) equation [2]. The key idea is to connect the time evolution of the value function to a probability density function evolving under an appropriate nonlinear Fokker–Planck equation and its mean-field representation [14, 10, 25, 11]. In this paper, we directly approximate the gradient of the value function and derive appropriate Hamiltonian evolution equations; thus extending mean-field formulations to the PMP.
@@
The remainder of the paper is structured as follows. The following Section 2 summarizes key mathematical results for finite horizon optimal control problems at the level of the underlying HJB and Fokker–Planck equations [2, 19]. The novel mean-field formulation of the PMP is provided in Section 3 starting from an Eulerian variational perspective in terms of the HJB equation. Following [27, 9], the arising action functional is subsequently transformed into a Lagrangian formulation with trajectories following mean-field Hamiltonian ODEs. The infinite horizon optimal control problem is treated in Section 4. Here the key result is a pair of forward mean-field ODEs, which asymptotically deliver the desired closed loop control law. A number of numerical approximation techniques are discussed in Section 5; including an approximation based on discrete Schrödinger bridges [16, 28] as outlined in Section 5.2. The mean-field formulation is tested for two infinite horizon discounted cost (3) optimal control problems in Section 6. The first example is provided by a controlled inverted pendulum [18] while the second example is provided by a controlled Lorenz-63 system [12].
@@ 2. Mathematical background
in $(X_{t},P_{t})$ subject to the boundary conditions $X_{0}=a$ and $P_{T}=\nabla_{x}f(X_{T})$ [21]. Formulation (7) offers significant computational savings compared to (6) and has been widely used especially in the context of nonlinear model predictive control (NMPC) [7, 23]. The PMP for ODEs has been extended to SDEs (1) giving rise to the, so called, stochastic PMP with (7) replaced by a pair of forward-backward SDEs [2].
@@
3. Pontryagin mean-field formulation
In this section, we develop a novel mean-field approach to the PMP starting from an Eulerian perspective in the spirit of ideal fluid dynamics [27]. Specifically, we combine the HJB equation (6) and a Liouville equation for the flow of densities $\rho_{t}(x)$ under a single variational principle. In order to do so, we assume that the initial states $X_{0}$ are distributed according to a given distribution $p_{0}(x)$. Following [27], we then transform the variational principle into a Lagrangian framework in terms of evolving positions $X_{t}\in\mathbb{R}^{d_{x}}$ and momenta $P_{t}\in\mathbb{R}^{d_{x}}$. These equations are Hamiltonian and provide the novel PMP formulation. While we consider the finite horizon optimal control problem (2) in this section; an extension to infinite horizon optimal control problems can be found in Section 4.
@@ 3.1. Eulerian formulation
(11a) $\mathcal{S}(v,\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}v_{t}+\nabla_{x}v_{t}\cdot b+c-\frac{1}{2}\|RG^{\rm T}\nabla_{x}v_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,+$
@@ 3.2. Lagrangian formulation
Let us now convert (11) into a corresponding Lagrangian formulation [27]. Let $X_{\tau}(a)$ denotes the solution of (13) at time $\tau=t$ with initial condition $X_{0}=a$. Here we have introduced $\tau$ as a new time variable in order to distinguish the two time derivatives arising from the Lagrangian (total time-derivative) and Eulerian (partial time-derivative) perspective. Furthermore, the initial conditions $a\sim p_{0}$ may be viewed as labels in the sense of Lagrangian fluid dynamics. We therefore consider functions $x(a)$, $p(x)$, $u(a)$, and $\beta(a)$ and note that
@@
where we have introduced the momentum variable
@@
as well as the fact that the total time derivative $\partial_{\tau}v_{\tau}(X_{\tau}(a))$ contributes boundary terms only. Furthermore,
@@
The equations of motion in $(X_{\tau},P_{\tau})$ arise now from the phase space action principle [4, 9] applied to
@@
(22b) $\qquad\qquad\beta(a)\cdot\nabla_{x}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right)+\frac{1}{2}\Sigma:D_{x}^{2}\phi(x(a)).$
@@
for each $a\in\mathbb{R}^{d_{x}}$. The boundary conditions are $X_{0}(a)=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$. The equations (29) are of mean-field type through $\phi_{t}(x)$, which is determined by the constraint (27).
@@
(30d) $\qquad\qquad\nabla_{x}\left\{c(X_{t}^{(i)})-\frac{1}{2}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\mathcal{L}_{\beta_{t}}(\phi_{t}(X_{t})-X_{t}\cdot P_{t})\right\}$
(30e) $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\nabla_{x}c(X_{t}^{(i)})+\mathcal{L}_{\beta_{t}}\nabla_{x}\phi_{t}(X_{t}),$
@@
which is the generator of a reversible diffusion process with invariant measure $\rho_{t}$ [19]. Since $\mathcal{L}_{\beta_{t}}^{\dagger}\rho_{t}=0$ for the dual operator $\mathcal{L}_{\beta_{t}}^{\dagger}$ [19], this is also the choice of $\beta_{t}$ which makes the variation with respect to $\phi_{t}$ vanish. We recall that this choice of $\beta_{t}$ leads to $\rho_{t}=\pi_{t}$ with $\pi_{t}$ the law of $X_{t}$ arising from the SDE (1).
@@
subject to $X_{0}=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$ solves the stochastic optimal control problem. We note that such a decoupling is possible even when $\Sigma$ is singular or even zero. This is in difference to the classical as well as the stochastic PMP [2].
@@
where $u_{t}^{\rm ref}(x)$ is a known reference control; if available. In the context of NMPC [23, 7], such a control could arise from the last receding horizon control window.
@@ 4. Discounted cost optimal control problems
Here we have scaled the value function by $e^{-\gamma t}$; i.e. $\tilde{v}_{t}=e^{-\gamma t}v_{t}$. After transferring to the Lagrangian framework, we undo the scaling in $\tilde{P}_{t}=\nabla_{x}\tilde{v}_{t}$ to obtain the action functional (41).
Since we are interested in equilibrium states only; i.e. $t\to\infty$, we can run both equations forward in time resulting in
@@
(44b) $\dot{P}_{t}$ $=-\gamma P_{t}+\nabla_{x}H(X_{t},P_{t},\beta_{t}).$
Then $(X_{\infty}(a),P_{\infty}(a))$ define the optimal control law (28) via the implied $\nabla_{x}\phi_{\infty}(x)$. In this setting, the choice (34) for $\beta_{t}$ emerges as the most natural one.
We note that such a joint forward integration is not possible for forward-backward SDE formulations of the PMP [2].

### v3

@@ Abstract.
This papers outlines a novel extension of the classical Pontryagin minimum (maximum) principle to stochastic optimal control problems. Contrary to the well-known stochastic Pontryagin minimum principle involving forward-backward stochastic differential equations, the proposed formulation is deterministic and of mean-field type. We denote it the McKean–Pontryagin minimum principle. The Hamiltonian structure of the proposed McKean–Pontryagin minimum principle is achieved via the introduction of an appropriate gauge variable. The gauge freedom can be used to decouple the forward and reverse time equations; hence simplifying the solution of the underlying boundary value problem. We also consider infinite horizon discounted cost optimal control problems. In this case, the mean-field formulation allows converting the computation of the desired optimal control law into solving a pair of forward mean-field ordinary differential equations. The McKean–Pontryagin minimum principle is tested numerically for a controlled inverted pendulum and a controlled Lorenz-63 system.
@@ 1. Introduction
Here $B_{t}$ denotes $d_{x}$-dimensional Brownian motion, $\Sigma\in\mathbb{R}^{d_{x}\times d_{x}}$ the symmetric positive definite diffusion matrix, $R\in\mathbb{R}^{d_{u}\times d_{u}}$ a symmetric positive definite weight matrix, $G(x)\in\mathbb{R}^{d_{x}\times d_{u}}$ the possibly position-dependent control matrix, $c(x)\geq 0$ the running cost, $f(x)\geq 0$ the terminal cost, and $\gamma>0$ the discount factor. See, for example, [5, 24] for more details. We also introduce the weighted norm $\|\cdot\|_{R}$ via
@@
The aim is to find the closed loop control law $U_{t}=u_{t}(X_{t})$ that minimizes the cost (2) or (3), respectively, over the set of admissible control laws.
@@
It is well known that the associated optimal control problems for ordinary differential equations (ODEs); i.e. $\Sigma=0$ in (1), can be solved by the Pontryagin minimum (maximum) principle (PMP) [28]. The PMP has been extended to SDEs (1) in terms of forward-backward SDEs [27, 5, 2]. In this paper, we provide an alternative extension via a deterministic mean-field formulation, which leads to interacting particle implementations in the form of Hamiltonian ODEs. The formulation is attractive from a computational perspective since it allows for a decoupling of the forward and backward evolution equations for finite horizon optimal control problems and a forward simulation for infinite horizon optimal control problems. Our mean-field approach delivers closed loop control laws in contrast to the classical PMP for ODEs [28]. The proposed methodology can then also be used to approximate viscosity solutions for controlled ODEs by setting $\Sigma=\epsilon I$ for sufficiently small $\epsilon>0$ [6].
@@
Mean-field formulations and their interacting particle approximations have been considered before in the context of the Hamilton–Jacobi–Bellman (HJB) equation [5, 2]. The key idea is to connect the time evolution of the value function to a probability density function evolving under an appropriate nonlinear Fokker–Planck equation and its mean-field representation [20, 16, 32, 17]. In this paper, we directly approximate the gradient of the value function and derive appropriate Hamiltonian evolution equations; thus extending mean-field formulations to the PMP. We denote our mean-field formulation the McKean–Pontryagin minimum principle.
As a further alternative, we mention the application of direct optimization methods to learn a parametric form of the desired control laws in the form of $u_{t}(x;\theta)$ for finite horizon cost functions (2). See, for example, [8, 9, 25, 15], which rely on deep learning approximation architectures. Similarly, infinite horizon discounted cost problems can also be reformulated as optimization problems in parameters $\theta$ based on approximations of the underlying stationary HJB equation; an approach pioneered in reinforcement learning [24].
@@
The remainder of the paper is structured as follows. The following Section 2 summarizes key mathematical results for finite horizon optimal control problems at the level of the underlying HJB and Fokker–Planck equations [5, 26]. The novel mean-field formulation of the PMP is provided in Section 3 starting from an Eulerian variational perspective in terms of the HJB equation. Following [34, 14], the arising action functional is subsequently transformed into a Lagrangian formulation with trajectories following mean-field Hamiltonian ODEs. The infinite horizon optimal control problem is treated in Section 4. Here the key result is a pair of forward mean-field ODEs, which asymptotically deliver the desired closed loop control law. A number of numerical approximation techniques are discussed in Section 6; including an approximation based on discrete Schrödinger bridges [22, 35] as outlined in Section 6.2. The McKean–Pontryagin formulation is tested for two infinite horizon discounted cost (3) optimal control problems in Section 7. The first example is provided by a controlled inverted pendulum [24] while the second example concerns a controlled Lorenz-63 system [18]. The final numerical example from Section 7.3 presents results from an implementation of the McKean–Pontryagin formulation in a receding horizon setting applied again to a controlled inverted pendulum.
@@ 2. Mathematical background
in $(X_{t},P_{t})$ subject to the boundary conditions $X_{0}=a$ and $P_{T}=\nabla_{x}f(X_{T})$ [28]. Formulation (7) offers significant computational savings compared to (6) and has been used especially in the context of nonlinear model predictive control (NMPC) [11, 30]. The PMP for ODEs has been extended to SDEs (1) giving rise to the, so called, stochastic PMP with (7) replaced by a pair of forward-backward SDEs [5, 2].
@@
3. McKean–Pontryagin Minimum Principle
In this section, we develop a novel mean-field approach to the PMP starting from an Eulerian perspective in the spirit of ideal fluid dynamics [34]. Specifically, we combine the HJB equation (6) and a Liouville equation for the flow of densities $\rho_{t}(x)$ under a single variational principle. In order to do so, we assume that the initial states $X_{0}=a$ in (1) are no longer fixed but distributed instead according to a given distribution $p_{0}(x)$. Following [34], we then transform the variational principle into a Lagrangian framework in terms of evolving positions $X_{t}\in\mathbb{R}^{d_{x}}$ and momenta (co-states) $P_{t}\in\mathbb{R}^{d_{x}}$. These equations are Hamiltonian and provide the novel McKean–Pontryagin formulation. While we consider the finite horizon optimal control problem (2) in this section; an extension to infinite horizon optimal control problems can be found in Section 4.
@@ 3.1. Eulerian formulation
(11a) $\mathcal{S}(v,\rho,\beta)$ $=-\int_{0}^{T}\int_{\mathbb{R}^{d_{x}}}\left\{\partial_{t}v_{t}+\nabla_{x}v_{t}\cdot b+c-\frac{1}{2}\|RG^{\rm T}\nabla_{x}v_{t}\|^{2}_{R}\right\}\rho_{t}{\rm d}x\,{\rm d}t\,-$
@@ 3.2. Lagrangian formulation
Let us convert (11) into a corresponding Lagrangian formulation [34]. Let $X_{\tau}(a)$ denotes the solution of (13) at time $\tau=t$ with initial condition $X_{0}=a$. Here we have introduced $\tau$ as a new time variable in order to distinguish the two time derivatives arising from the Lagrangian (total time-derivative) and Eulerian (partial time-derivative) perspective. Furthermore, the initial conditions $a\sim p_{0}$ may be viewed as labels in the sense of Lagrangian fluid dynamics. We therefore consider functions $x(a)$, $p(x)$, $u(a)$, and $\beta(a)$ and note that
@@
where we have introduced the momentum variable (co-state)
@@
as well as the fact that the total time derivative $\partial_{\tau}v_{\tau}(X_{\tau}(a))$ contributes boundary terms only to action functional (17). Furthermore,
@@
The equations of motion in $(X_{\tau},P_{\tau})$ arise now from the Hamiltonian phase space action principle [7, 14] applied to
@@
(22b) $\qquad\qquad\beta(a)^{\rm T}\nabla_{x}\left(\phi(x(a))-x(a)^{\rm T}p(a)\right)+\frac{1}{2}\Sigma:D_{x}^{2}\phi(x(a)).$
@@
for each $a\in\mathbb{R}^{d_{x}}$. The boundary conditions are $X_{0}(a)=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$. The equations (29) are of mean-field type through $\phi_{\tau}(x)$, which is determined by the constraint (27).
@@
(30d) $\qquad\qquad\nabla_{x}\left\{c(X_{t})-\frac{1}{2}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\mathcal{L}_{\beta_{t}}(\phi_{t}(X_{t})-X_{t}\cdot P_{t})\right\}$
(30e) $=\left(D_{x}b(X_{t})\right)^{\rm T}P_{t}-\frac{1}{2}\nabla_{x}\|RG(X_{t})^{\rm T}P_{t}\|^{2}_{R}+\nabla_{x}c(X_{t})+\mathcal{L}_{\beta_{t}}\nabla_{x}\phi_{t}(X_{t}),$
@@
which is the generator of a reversible diffusion process with invariant measure $\rho_{t}$ [26]. Since $\mathcal{L}_{\beta_{t}}^{\dagger}\rho_{t}=0$ for the dual operator $\mathcal{L}_{\beta_{t}}^{\dagger}$ [26], this is also the choice of $\beta_{t}$ which makes the variation with respect to $\phi_{t}$ vanish. We recall that (15) leads to $\rho_{t}=\pi_{t}$ with $\pi_{t}$ the law of $X_{t}$ arising from the SDE (1).
@@
subject to $X_{0}=a\sim p_{0}$ and $P_{T}(a)=\nabla_{x}f(X_{T}(a))$ solves the stochastic optimal control problem. We note that such a decoupling is possible even when $\Sigma$ is singular or indeed zero. This is in difference to the classical as well as the stochastic PMP [5].
@@
where $u_{t}^{\rm ref}(x)$ is a known reference control; if available. In the context of NMPC [30, 11], such a control could arise from the previous receding horizon control window.
@@ 4. Discounted cost optimal control problems
Here we have scaled the value function by $e^{-\gamma t}$; i.e. $\tilde{v}_{t}=e^{-\gamma t}v_{t}$. After switching to the Lagrangian framework, we undo the scaling in $\tilde{P}_{t}=\nabla_{x}\tilde{v}_{t}$ to obtain the action functional (41). The associated equations of motion are
@@
(44b) $-\dot{P}_{t}$ $=-\gamma P_{t}+\nabla_{x}H(X_{t},P_{t},\beta_{t}).$
The evolution equations (44) are still posed as a boundary value problem with (44a) solved forward in time and (44b) backward in time, respectively.
We call a solution of (44) with $\partial_{t}\phi_{t}=0$ a steady state solution and denote it by $\hat{\phi}$. Following (42), $\hat{\phi}$ satisfies the stationary HJB equation
(45) $0=-\gamma\hat{\phi}+b\cdot\nabla_{x}\hat{\phi}+\frac{1}{2}\nabla_{x}\cdot(\Sigma\nabla_{x}\hat{\phi})+c-\frac{1}{2}\|RG(x)^{\rm T}\nabla_{x}\hat{\phi}\|^{2}_{R}.$
Note that the positions $X_{t}$ can still undergo a time evolution; but need to satisfy
(46) $\dot{P}_{t}=D_{x}^{2}\hat{\phi}(X_{t})\dot{X}_{t},$
which follows from (27). If we are interested in approximating such steady state solutions, we can run both equations forward in time by replacing $\dot{P}_{t}$ by $-\dot{P}_{t}$ in (44) and adding an appropriate advection term, which results in
(47a) $\dot{X}_{t}$ $=\nabla_{p}H(X_{t},\Pi_{t},\beta_{t}),$
(47b) $\dot{\Pi}_{t}$ $=-\gamma\Pi_{t}+\nabla_{x}H(X_{t},\Pi_{t},\beta_{t})+2D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}.$
Here we have introduced new momenta (co-states) $\Pi_{t}$ for clarity of presentation.
Theorem 4.1.
Let us assume that $(X_{t},P_{t})$, $t\in\mathbb{R}$ represent a steady state solution of (44); i.e., (46) holds for suitable $\hat{\phi}$. Then solutions $(X_{t},\Pi_{t})$ of (47), $t\geq 0$, with initial condition $(X_{0},\Pi_{0})=(X_{0},P_{0})$ also represent a steady state solution.
Proof.
We recall the definition $\tilde{y}_{t}(x)=\nabla_{t}\phi_{t}(x)$ and follow the proof of Theorem 3.1. We find that (47) implies
(48) $\dot{\Pi}_{t}=-\partial_{t}\tilde{y}_{t}(X_{t})+D_{x}^{2}\phi_{t}(X_{t})\dot{X}_{t}$
with $\partial_{t}\tilde{y}_{t}$ now satisfying
(49) $\partial_{t}\tilde{y}_{t}=-\gamma\tilde{y}_{t}+\nabla_{x}(\tilde{y}_{t}\cdot b)+\frac{1}{2}\nabla_{x}(\Sigma:D_{x}\tilde{y}_{t})+\nabla_{x}c-\frac{1}{2}\nabla_{x}\|RG^{\rm T}\tilde{y}_{t}\|^{2}_{R}$
instead of (33). It follows from $\phi_{t}=\hat{\phi}$ that
(50) $\dot{\Pi}_{t}=D_{x}^{2}\hat{\phi}(X_{t})\dot{X}_{t}$
in agreement with (46). In other words, $(X_{t},P_{t})=(X_{t},\Pi_{t})$ for all $t\geq 0$ when initialized in steady state and steady state solutions are preserved under (47). ∎
We close this section by mentioning that (47) is related to iterative approaches for approximating the stationary value function (or $Q$-function) used in reinforcement learning and which are based on the stationary Bellman equation [24]. Our approach targets instead the gradients of the value function (momenta/co-states) directly. This has been found beneficial from a theoretical perspective [5] and provides flexibility in terms of constraints added on the states and controls. It is also often found that approximating the gradient directly instead of its underlying function is more robust computationally. Compare, for example, density versus score estimation [21, 13].
5. Linear control problems
We analyze the proposed formulations in the setting of controlled linear SDEs
(51) ${\rm d}X_{t}=AX_{t}+GU_{t}+\Sigma^{1/2}{\rm d}B_{t},\qquad X_{0}\sim{\rm N}(0,C_{0})$
for infinite time discounted cost functional (3) with running cost
(52) $c(x)=\frac{1}{2}x^{\rm T}Cx$
and $R=I$. We write the function $\phi_{t}(x)$ as
(53) $\phi_{t}(x)=\frac{1}{2}x^{\rm T}\Omega_{t}x.$
With this ansatz the Hamiltonian equations of motion (29) take the form
(54a) $\dot{X}_{t}$ $=AX_{t}-GG^{\rm T}P_{t}-\beta_{t},$
(54b) $-\dot{P}_{t}$ $=-\gamma P_{t}+A^{\rm T}P_{t}+CX_{t}+\Omega_{t}\beta_{t}.$
The choice (34) for the gauge variable $\beta_{t}$ leads to
(55) $\beta_{t}=-\frac{1}{2}\Sigma C_{t}^{-1}X_{t}$
with $C_{t}$ the covariance matrix of the Gaussian distributed $X_{t}$ with mean zero. The constraint (27) becomes
(56) $P_{t}=\Omega_{t}X_{t},$
which yields
(57a) $-\dot{\Omega}_{t}X_{t}$ $=\Omega_{t}\dot{X}_{t}-\dot{P}_{t}$
(57b) $=\Omega_{t}\left(AX_{t}-GG^{\rm T}P_{t}-\beta_{t}\right)-\gamma P_{t}+A^{\rm T}P_{t}+CX_{t}+\Omega_{t}\beta_{t}$
(57c) $=\left(-\gamma\Omega_{t}+\Omega_{t}A-\Omega_{t}GG^{\rm T}\Omega_{t}+A^{\rm T}\Omega_{t}+C\right)X_{t}$
and from which we can read off the backward Riccati equation for $\Omega_{t}$ [5]:
(58) $-\dot{\Omega}=-\gamma\Omega_{t}+\Omega_{t}A+A^{\rm T}\Omega_{t}-\Omega_{t}GG^{\rm T}\Omega_{t}+C.$
Let us now investigate the forward evolution equations (47). We obtain
(59a) $\dot{X}_{t}$ $=AX_{t}-GG^{\rm T}P_{t}-\beta_{t}$
(59b) $\dot{\Pi}_{t}$ $=-\gamma\Pi_{t}+A^{\rm T}\Pi_{t}+CX_{t}+\Omega_{t}\beta_{t}+2\Omega_{t}\dot{X}_{t}$
subject to the constraint $\Pi_{t}=\Omega_{t}X_{t}$. This time we obtain
(60a) $\dot{\Omega}_{t}X_{t}$ $=-\Omega_{t}\dot{X}_{t}+\dot{\Pi}_{t}$
(60b) $=\Omega_{t}\left(AX_{t}-GG^{\rm T}\Pi_{t}-\beta_{t}\right)-\gamma\Pi_{t}+A^{\rm T}\Pi_{t}+CX_{t}+\Omega_{t}\beta_{t}$
(60c) $=\left(-\gamma\Omega_{t}+\Omega_{t}A+A^{\rm T}\Omega_{t}-\Omega_{t}GG^{\rm T}\Omega_{t}+C\right)X_{t}$
and read off the forward Riccati equation
(61) $\dot{\Omega}=-\gamma\Omega_{t}+\Omega_{t}A+A^{\rm T}\Omega_{t}-\Omega_{t}GG^{\rm T}\Omega_{t}+C.$
Clearly, the forward Riccati equation is stable if and only if the backward Riccati equation is stable and their stationary solutions agree. Of course, our approach is to propagate $X_{t}$ and $\Pi_{t}$ and to then find the appropriate $\Omega_{t}$ by enforcing the constraint $\Pi_{t}=\Omega_{t}X_{t}$ via regression; i.e.
(62) $\Omega_{t}=\arg\min_{\Omega\in\mathbb{R}^{d_{x}\times d_{x}}}\mathbb{E}\left[\|\Pi_{t}-\Omega X_{t}\|^{2}\right].$
We describe numerical implementations for the general nonlinear setting in the subsequent section.